# Symmetric Iterations with Countable Support: A Framework for Choiceless ZF Extensions

### v1

@@ 3.3 Hereditarily Symmetric Names in Countable Support Iterations
Having constructed the filter $\tilde{\mathcal{F}}_{\lambda}$ at limit stages—via block factorization when $\mathrm{cf}(\lambda)=\omega$ (inverse limit) and stage-bounding when $\mathrm{cf}(\lambda)\geq\omega_{1}$ (direct limit)—and established its $\omega_{1}$-completeness (Section 3.2, Theorem 3.15), we now define the class of hereditarily symmetric names and verify its basic closure properties in the countable-support setting.
Definition 3.16 (Hereditarily Symmetric Names — Iterated, Countable Support Case).
A $\mathbb{P}_{\lambda}$-name $\dot{x}$ is hereditarily symmetric (HS) (with respect to $\tilde{\mathcal{F}}_{\lambda}$) if:
  1. 1.
    $\mathrm{fix}_{\mathcal{G}_{\lambda}}(\dot{x})\in\tilde{\mathcal{F}}_{\lambda}$, and
  2. 2.
    for every $(\dot{y},p)\in\dot{x}$, the name $\dot{y}$ is HS.
$\mathrm{fix}\!\bigl(\langle\dot{x}_{j}:j\in J\rangle\bigr)\;=\;\bigcap_{j\in J}\mathrm{fix}(\dot{x}_{j}).$ (1)
@@
Let $\dot{s}=\langle\dot{x}_{j}:j\in J\rangle$ be the canonical sequence name. Then $g\in\mathrm{fix}(\dot{s})$ iff $g(\dot{s})=\dot{s}$, which (by the recursive definition of the action on names) holds iff $g(\dot{x}_{j})=\dot{x}_{j}$ for all $j\in J$, i.e. iff $g\in\bigcap_{j\in J}\mathrm{fix}(\dot{x}_{j})$. ∎
@@
The class of HS names is closed under pairing, union, formation of subsets by bounded formulas (Separation instances over a fixed ambient set), and the usual forcing operations (e.g. check names).
@@
$\mathrm{fix}\!\bigl(\langle\dot{x},\dot{y}\rangle\bigr)\;=\;\mathrm{fix}(\dot{x})\cap\mathrm{fix}(\dot{y}),$
@@
$\mathrm{fix}(\dot{s})\ =\ \bigcap_{n<\omega}\mathrm{fix}(\dot{x}_{n}).$
Since each $\mathrm{fix}(\dot{x}_{n})\in\tilde{\mathcal{F}}_{\lambda}$ and $\tilde{\mathcal{F}}_{\lambda}$ is $\omega_{1}$–complete (Theorem 3.15), we have $\mathrm{fix}(\dot{s})\in\tilde{\mathcal{F}}_{\lambda}$, so $\dot{s}$ is HS.
@@
If $g\in\mathrm{fix}(\dot{A})$ then $g(\dot{U})=\dot{U}$ (because $g$ just transports witnesses in the defining formula), so $\mathrm{fix}(\dot{U})\supseteq\mathrm{fix}(\dot{A})\in\tilde{\mathcal{F}}_{\lambda}$. Hence $\dot{U}$ is HS.
Bounded Separation (over a fixed ambient set). Let $\dot{B}=\{z\in\dot{A}:\varphi(z,\bar{u})\}$ where $\dot{A}$ is HS and $\bar{u}$ is a tuple of HS parameters, and $\varphi$ is a bounded formula. If $g$ fixes $\dot{A}$ and $\bar{u}$, then by automorphism–invariance of truth for bounded formulas in the forcing language, $g$ fixes $\dot{B}$. Hence
$\mathrm{fix}(\dot{B})\ \supseteq\ \mathrm{fix}(\dot{A})\cap\mathrm{fix}(\bar{u}),$
@@
Check names and standard operations. If $a\in V$, then $\check{a}$ is fixed by all $g\in\mathcal{G}_{\lambda}$, so $\mathrm{fix}(\check{a})=\mathcal{G}_{\lambda}\in\tilde{\mathcal{F}}_{\lambda}$. Other usual name constructors are handled similarly; when a constructor uses only finitely many HS inputs, we use finite-intersection closure of $\tilde{\mathcal{F}}_{\lambda}$, and when it packages countably many HS inputs (e.g. countable tuples) we use (1) together with $\omega_{1}$–completeness (Theorem 3.15).
@@
Remark 3.18 (Ablation checkpoints for Theorem 3.17).
  1. (a)
    Drop $\omega_{1}$–completeness: Replacement’s final stabilizer intersection may leave the filter.
  2. (b)
    Drop normality: Separation can fail (stabilizers of parameters + set need not generate a filter element).
  3. (c)
    Drop support bounds: the tuple–stabilizer reduction no longer yields a generator.
@@ 3.4 $\mathrm{ZF}$ Preservation
Write $\mathrm{fix}(\dot{x})=\mathrm{fix}_{\mathcal{G}_{\lambda}}(\dot{x})$ and let $\tilde{\mathcal{F}}_{\lambda}$ be the normal, $\omega_{1}$–complete limit filter from Theorem 3.15 (constructed at $\lambda$ as in Definition 3.11). We also use the tuple–stabilizer identity (1).
@@
Pairing and Union. If $\dot{x},\dot{y}$ are HS, then by (1) $\mathrm{fix}(\langle\dot{x},\dot{y}\rangle)=\mathrm{fix}(\dot{x})\cap\mathrm{fix}(\dot{y})\in\tilde{\mathcal{F}}_{\lambda}$, so the ordered pair is HS; hereditariness gives $\{\dot{x},\dot{y}\}$ HS as well.
@@
If $g\in\mathrm{fix}(\dot{A})$ then $g(\dot{U})=\dot{U}$ (since $g$ transports witnesses in the defining formula), so $\mathrm{fix}(\dot{U})\supseteq\mathrm{fix}(\dot{A})\in\tilde{\mathcal{F}}_{\lambda}$. Hence $\dot{U}$ is HS and therefore $U:=\dot{U}^{G}=\bigcup(\dot{A}^{G})$ belongs to $M$.
@@
If $g\in\mathrm{fix}(\dot{A})\cap\mathrm{fix}(\vec{\dot{p}})$, then by automorphism invariance of the forcing relation, $p\Vdash\varphi(\sigma,\vec{\dot{p}})\iff g(p)\Vdash\varphi(g\sigma,\vec{\dot{p}})$, whence $g\in\mathrm{fix}(\dot{B})$. Thus $\mathrm{fix}(\dot{B})\supseteq\mathrm{fix}(\dot{A})\cap\mathrm{fix}(\vec{\dot{p}})\in\tilde{\mathcal{F}}_{\lambda}$, so $\dot{B}$ is HS and $B:=\dot{B}^{G}=\{x\in A:\varphi^{M}(x,\vec{p})\}$. Hence $M$ satisfies Separation.
@@
Then $\mathrm{HS}(\dot{a})$ is a set in $V$ (it is a definable subset of $\mathrm{Nm}_{\leq\alpha}$), and if $g\in\mathrm{fix}(\dot{a})$ then $g$ preserves both subsetting $\dot{a}$ and the HS property, so $\mathrm{HS}(\dot{a})$ is invariant under $\mathcal{G}_{\lambda}$. Consequently,
@@
If $g\in\mathrm{fix}(\dot{A})\cap\mathrm{fix}(\dot{f})$, then by definition $g\in\mathrm{fix}(\dot{\mathrm{rng}})$; hence $\mathrm{fix}(\dot{\mathrm{rng}})\supseteq\mathrm{fix}(\dot{A})\cap\mathrm{fix}(\dot{f})\in\tilde{\mathcal{F}}_{\lambda}$, so $\dot{\mathrm{rng}}$ is HS. By the forcing theorem and functionality of $\dot{f}$, we have
@@
In the countable-support setting, $\omega_{1}$-completeness of $\tilde{\mathcal{F}}_{\lambda}$ is needed to close HS under countable intersections of stabilizers (e.g., for closure under countable unions of HS constituents). If one proves Replacement by first enumerating the domain and setting $\dot{B}=\{\dot{y}_{n}:n<\omega\}$, then $\omega_{1}$-completeness is used to ensure $\bigcap_{n<\omega}\mathrm{fix}(\dot{y}_{n})\in\tilde{\mathcal{F}}_{\lambda}$.
In the proof of Theorem 3.20 given above, we instead use the relational range name, so the Replacement step itself only requires the finite intersection $\mathrm{fix}(\dot{A})\cap\mathrm{fix}(\dot{f})\in\tilde{\mathcal{F}}_{\lambda}$. Thus, for countable support the real novelty manifests earlier (in closure under countable unions and similar uses of $\omega_{1}$-completeness), rather than in the Replacement step per se.
@@ 3.5 Dependent Choice and Choice Fragments
Definition 3.22 (Dependent Choice (DC)).
For a nonempty set $A$ and relation $R\subseteq A\times A$ with $\forall x\in A\,\exists y\in A\ (xRy)$, DC asserts the existence of a sequence $(a_{n})_{n<\omega}$ with $a_{n}Ra_{n+1}$ for all $n<\omega$.
Theorem 3.23 (DC Preservation).
Assume the ambient ground satisfies $\mathrm{ZFC}$. If $\tilde{\mathcal{F}}_{\lambda}$ is $\omega_{1}$-complete, then $M=V(\mathbb{P}_{\lambda})_{\mathrm{HS}}\models\mathrm{DC}$.
@@
Since $M$ is transitive (Theorem 3.20), the same statement holds in $V[G]$. But $V[G]\models\mathrm{ZFC}$, hence $V[G]\models\mathrm{DC}$, so there exists a sequence $f:\omega\to A$ in $V[G]$ such that $R(f(n),f(n+1))$ for all $n$. Since $f$ is an $\omega$-sequence of elements of $M$, it suffices to show $M$ is closed under countable sequences.
@@
Remark 3.24 (Scope of DC in the countable-support setting).
In the countable-support case, $\omega_{1}$-completeness of the limit filter yields $DC_{<\omega_{1}}$; in particular $DC=DC_{\omega}$ holds in the symmetric model.
@@
3.6 Iteration length and class-length versions

### v3

@@ 3.3 Hereditarily Symmetric Names in Countable Support Iterations
Having defined the limit filter $\tilde{\mathcal{F}}_{\lambda}$ (Definition 3.9) and verified its normality and $\omega_{1}$–completeness (Theorem 3.13), we now define the class of hereditarily symmetric names at stage $\lambda$ and verify the standard closure properties needed later for $\mathrm{ZF}$ and $\mathrm{DC}$ preservation.
Definition 3.14 (HS at stage $\lambda$).
Let $\langle\mathbb{P}_{\lambda},\mathcal{G}_{\lambda},\widetilde{\mathcal{F}}_{\lambda}\rangle$ be the symmetric system at stage $\lambda$. We write
$\mathrm{HS}_{\lambda}:=\mathrm{HS}(\mathbb{P}_{\lambda},\mathcal{G}_{\lambda},\widetilde{\mathcal{F}}_{\lambda})$
for the class of hereditarily $\widetilde{\mathcal{F}}_{\lambda}$-symmetric $\mathbb{P}_{\lambda}$-names, as in Definition 2.10. When $\lambda$ is fixed, we simply write $\mathrm{HS}$.
$\operatorname{sym}\!\bigl(\langle\dot{x}_{j}:j\in J\rangle\bigr)\;=\;\bigcap_{j\in J}\operatorname{sym}(\dot{x}_{j}).$ (1)
@@
Let $\dot{s}=\langle\dot{x}_{j}:j\in J\rangle$ be the canonical sequence name. Then $g\in\operatorname{sym}(\dot{s})$ iff $g(\dot{s})=\dot{s}$, which (by the recursive definition of the action on names) holds iff $g(\dot{x}_{j})=\dot{x}_{j}$ for all $j\in J$, i.e. iff $g\in\bigcap_{j\in J}\operatorname{sym}(\dot{x}_{j})$. ∎
@@
The class of HS names is closed under pairing, countable tuples, union, formation of subsets by bounded formulas (Separation instances over a fixed ambient set), and the usual forcing operations (e.g. check names).
@@
$\operatorname{sym}\!\bigl(\langle\dot{x},\dot{y}\rangle\bigr)\;=\;\operatorname{sym}(\dot{x})\cap\operatorname{sym}(\dot{y}),$
@@
$\operatorname{sym}(\dot{s})\ =\ \bigcap_{n<\omega}\operatorname{sym}(\dot{x}_{n}).$
Since each $\operatorname{sym}(\dot{x}_{n})\in\tilde{\mathcal{F}}_{\lambda}$ and $\tilde{\mathcal{F}}_{\lambda}$ is $\omega_{1}$–complete (Theorem 3.13), we have $\operatorname{sym}(\dot{s})\in\tilde{\mathcal{F}}_{\lambda}$, so $\dot{s}$ is HS.
@@
If $g\in\operatorname{sym}(\dot{A})$ then $g(\dot{U})=\dot{U}$ (because $g$ just transports witnesses in the defining formula), so $\operatorname{sym}(\dot{U})\supseteq\operatorname{sym}(\dot{A})\in\tilde{\mathcal{F}}_{\lambda}$. Hence $\dot{U}$ is HS.
Bounded Separation (over a fixed ambient set). Let $\dot{B}=\{z\in\dot{A}:\varphi(z,\bar{u})\}$ where $\dot{A}$ is HS and $\bar{u}$ is a tuple of HS parameters, and $\varphi$ is a bounded formula. If $g$ fixes $\dot{A}$ and $\bar{u}$, then by the Symmetry Lemma (Lemma 2.9), $g$ fixes $\dot{B}$. Hence
$\operatorname{sym}(\dot{B})\ \supseteq\ \operatorname{sym}(\dot{A})\cap\operatorname{sym}(\bar{u}),$
@@
Check names and standard operations. If $a\in V$, then $\check{a}$ is fixed by all $g\in\mathcal{G}_{\lambda}$, so $\operatorname{sym}(\check{a})=\mathcal{G}_{\lambda}\in\tilde{\mathcal{F}}_{\lambda}$. Other usual name constructors are handled similarly; when a constructor uses only finitely many HS inputs, we use finite-intersection closure of $\tilde{\mathcal{F}}_{\lambda}$, and when it packages countably many HS inputs (e.g. countable tuples) we use (1) together with $\omega_{1}$–completeness (Theorem 3.13).
@@ 3.4 $\mathrm{ZF}$ Preservation
Write $\operatorname{sym}(\dot{x})=\operatorname{sym}_{\mathcal{G}_{\lambda}}(\dot{x})$ and let $\tilde{\mathcal{F}}_{\lambda}$ be the normal, $\omega_{1}$–complete limit filter from Theorem 3.13 (constructed at $\lambda$ as in Definition 3.9). We also use the tuple–stabilizer identity (1).
@@
Pairing and Union. If $\dot{x},\dot{y}$ are HS, then by (1) $\operatorname{sym}(\langle\dot{x},\dot{y}\rangle)=\operatorname{sym}(\dot{x})\cap\operatorname{sym}(\dot{y})\in\tilde{\mathcal{F}}_{\lambda}$, so the ordered pair is HS; hereditariness gives $\{\dot{x},\dot{y}\}$ HS as well.
@@
If $g\in\operatorname{sym}(\dot{A})$ then $g(\dot{U})=\dot{U}$ (since $g$ transports witnesses in the defining formula), so $\operatorname{sym}(\dot{U})\supseteq\operatorname{sym}(\dot{A})\in\tilde{\mathcal{F}}_{\lambda}$. Hence $\dot{U}$ is HS and therefore $U:=\dot{U}^{G}=\bigcup(\dot{A}^{G})$ belongs to $M$.
@@
If $g\in\operatorname{sym}(\dot{A})\cap\operatorname{sym}(\vec{\dot{p}})$, then by the Symmetry Lemma (Lemma 2.9), $p\Vdash\varphi(\sigma,\vec{\dot{p}})\iff g(p)\Vdash\varphi(g\sigma,\vec{\dot{p}})$, whence $g\in\operatorname{sym}(\dot{B})$. Thus $\operatorname{sym}(\dot{B})\supseteq\operatorname{sym}(\dot{A})\cap\operatorname{sym}(\vec{\dot{p}})\in\tilde{\mathcal{F}}_{\lambda}$, so $\dot{B}$ is HS and $B:=\dot{B}^{G}=\{x\in A:\varphi^{M}(x,\vec{p})\}$. Hence $M$ satisfies Separation.
@@
Then $\mathrm{HS}(\dot{a})$ is a set in $V$ (it is a definable subset of $\mathrm{Nm}_{\leq\alpha}$), and if $g\in\operatorname{sym}(\dot{a})$ then $g$ preserves both subsetting $\dot{a}$ and the HS property, so $\mathrm{HS}(\dot{a})$ is invariant under $\mathcal{G}_{\lambda}$. Consequently,
@@
If $g\in\operatorname{sym}(\dot{A})\cap\operatorname{sym}(\dot{f})$, then by definition $g\in\operatorname{sym}(\dot{\mathrm{rng}})$; hence $\operatorname{sym}(\dot{\mathrm{rng}})\supseteq\operatorname{sym}(\dot{A})\cap\operatorname{sym}(\dot{f})\in\tilde{\mathcal{F}}_{\lambda}$, so $\dot{\mathrm{rng}}$ is HS. By the forcing theorem and functionality of $\dot{f}$, we have
@@
In the countable-support setting, $\omega_{1}$-completeness of $\tilde{\mathcal{F}}_{\lambda}$ is needed to close HS under countable intersections of stabilizers (e.g., for closure under countable unions of HS constituents). If one proves Replacement by first enumerating the domain and setting $\dot{B}=\{\dot{y}_{n}:n<\omega\}$, then $\omega_{1}$-completeness is used to ensure $\bigcap_{n<\omega}\operatorname{sym}(\dot{y}_{n})\in\tilde{\mathcal{F}}_{\lambda}$.
In the proof of Theorem 3.17 given above, we instead use the relational range name, so the Replacement step itself only requires the finite intersection $\operatorname{sym}(\dot{A})\cap\operatorname{sym}(\dot{f})\in\tilde{\mathcal{F}}_{\lambda}$. Thus, for countable support the real novelty manifests earlier (in closure under countable unions and similar uses of $\omega_{1}$-completeness), rather than in the Replacement step per se.
@@ 3.5 Dependent Choice and Choice Fragments
Remark 3.19.
We use $\mathrm{DC}$ in the sense of Definition 2.11 (the set formulation).
Theorem 3.20 (DC Preservation (set-length stages)).
Assume $\lambda$ is a set ordinal and the ambient ground satisfies $\mathrm{ZFC}$. If $\tilde{\mathcal{F}}_{\lambda}$ is $\omega_{1}$-complete, then $M=V(\mathbb{P}_{\lambda})_{\mathrm{HS}}\models\mathrm{DC}$.
@@
Since $M$ is transitive (Theorem 3.17), the same statement holds in $V[G]$. Since $\mathbb{P}_{\lambda}$ is set forcing and $V\models\mathrm{ZFC}$, the forcing extension $V[G]\models\mathrm{ZFC}$, hence $V[G]\models\mathrm{DC}$. Therefore there exists a sequence $f:\omega\to A$ in $V[G]$ such that $R(f(n),f(n+1))$ for all $n$. Since $f$ is an $\omega$-sequence of elements of $M$, it suffices to show $M$ is closed under countable sequences.
@@
Remark 3.21 (Scope of DC obtained here).
The argument of Theorem 3.20 yields $\mathrm{DC}=\mathrm{DC}_{\omega}$ (the set formulation from Definition 2.11). Related preservation results for stronger fragments of dependent choice under $\kappa$-complete symmetry filters (under additional hypotheses on the forcing) appear, e.g., in [4, 1]. We do not claim any uncountable-length dependent choice (e.g. $\mathrm{DC}_{\omega_{1}}$) without additional hypotheses.
Corollary 3.22 (No infinite Dedekind-finite sets; no amorphous sets).
Under the hypotheses of Theorem 3.20, the symmetric model $M$ contains no infinite Dedekind-finite set and, in particular, no amorphous set.
Proof.
It is a theorem of $\mathrm{ZF}+\mathrm{DC}$ that every infinite set contains a countably infinite subset (see e.g. [2, Theorem 8.2]). Since $M\models\mathrm{DC}$ by Theorem 3.20, this applies inside $M$. An amorphous set—an infinite set with no infinite proper subset—would in particular be an infinite set with no countably infinite subset, contradicting $\mathrm{DC}$ in $M$. ∎
Remark 3.23 (Contrast with finite-support symmetric iterations).
In finite-support symmetric iterations, $\mathrm{DC}$ is not preserved in general at limit stages of cofinality $\omega$ (Remark 3.25), and amorphous or Dedekind-finite sets can appear in the resulting symmetric model. The classical example is Cohen’s first model: forcing with $\mathrm{Add}(\omega,\omega)$ and taking the symmetric submodel under finite-support permutations of coordinates yields a set of Cohen reals $A=\{a_{n}:n<\omega\}$ that is infinite but Dedekind-finite in the symmetric model, because no $\omega$-sequence enumerating $A$ is hereditarily symmetric (the enumeration stabilizer requires fixing all coordinates, falling outside any finite-support filter). See e.g. [2, Chapter 15] for details.
In contrast, the present countable-support framework ensures $\omega_{1}$-completeness of the limit filter at every cofinality-$\omega$ stage (Theorem 3.13), which forces HS closure under $\omega$-tuples (Theorem 3.15). The “forget the enumeration” mechanism of the finite-support setting is therefore unavailable: any countable enumeration of an infinite set is itself an HS name, and hence belongs to $M$.
@@
Remark 3.25 (Why finite support does not suffice here).
(1) Countable sequences in the symmetric model. The proof of $\mathrm{DC}$ preservation (Theorem 3.20) does not rely on the forcing partial order being $\omega$-closed. Instead, the proof relies on $\omega_{1}$–completeness of the limit filter together with closure of $\mathrm{HS}$ under countable tuples (Theorem 3.15). Finite-support limit filters typically lack countable-intersection closure at limits of cofinality $\omega$, and therefore do not secure the tuple closure needed for $\mathrm{DC}$.
(2) Closure of symmetry under countable tuples. Our arguments repeatedly use that if $\langle\dot{x}_{n}:n<\omega\rangle$ consists of HS names, then the tuple itself is HS, which requires
$\bigcap_{n<\omega}\operatorname{sym}(\dot{x}_{n})\in\tilde{\mathcal{F}}_{\lambda}.$
This is exactly the $\omega_{1}$-completeness provided by Theorem 3.13.
(3) Applications requiring $\omega_{1}$-complete automorphism groups at limit stages. In the iterated unordered pairs model (Section 4), the impossibility of a choice function for $\mathcal{F}$ rests on each individual swap $\sigma_{\alpha}$ not belonging to the stabilizer of any choice-function name. For a DC-sequence through the pairs, by contrast, the stabilizer of the sequence name needs only to contain $\bigcap_{n<\omega}\operatorname{sym}(\dot{x}_{n})$, which $\omega_{1}$-completeness of the limit filter guarantees (Theorem 3.13). No individual $\sigma_{n}$ need fix the sequence name; $\omega_{1}$-completeness is what allows the tuple to be hereditarily symmetric as a whole. In a finite-support iteration the limit filter at a cofinality-$\omega$ stage is closed only under finite intersections of head pullbacks, which is insufficient: as shown explicitly in Proposition 4.4, a DC-counterexample exists in the finite-support model at stage $\omega$, and no HS name for the required sequence can be constructed. $\mathrm{DC}$ preservation is therefore an intrinsic consequence of $\omega_{1}$-completeness, not a byproduct of some other construction.
3.6 Iteration length